# mode = title

### v1

@@ 6.2 Monocular Depth Estimation Module (MDEM)
Figure 6, 7 present the performances of the monocular depth estimation module, BodySLAM, compared to EndoDepth and EndoSfmLearner respectively on the Hamlyn and SCARED datasets.
@@
Figure 6: Comparison of depth estimation performance for BodySLAM, EndoSfmLearner, and EndoDepth across various metrics on Hamlyn Dataset. Box plots show that BodySLAM achieves the best performance with the lowest median values for absolute relative difference, RMSE, RMSE log, and squared relative difference, and the highest median values for accuracy thresholds $\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$. EndoSfmLearner demonstrates the second best performance, while EndoDepth shows the poorest results among the three methods compared.
@@
Figure 7: Comparison of depth estimation performance for BodySLAM, EndoSfmLearner, and EndoDepth across various metrics on Hamlyn Dataset. Box plots show that BodySLAM achieves the best performance with the lowest median values for absolute relative difference, RMSE, RMSE log, and squared relative difference, and the highest median values for accuracy thresholds $\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$. EndoSfmLearner demonstrates the second best performance, while EndoDepth shows the poorest results among the three methods compared. Outliers beyond 1.5 times the interquartile range from the quartiles have been excluded to enhance clarity and focus on the central trends of the data.
@@
The figures compare the three methods across multiple evaluation metrics for each dataset. The x-axis represents the different methods (BodySLAM, EndoDepth, and EndoSfmLearner), while the y-axis shows the values for the evaluation metrics. In the case of the Hamlyn dataset (Figure 6), BodySLAM outperforms both EndoDepth and EndoSfmLearner across all metrics. For Absolute relative difference, RMSE, RMSE log and Squared Relative error (first row, Figure 6), BodySLAM exhibits lower median values and smaller interquartile ranges compared to the other methods. Regarding the threshold accuracy (second row, Figure 6), where values close to or equal to 1 are considered optimal, BodySLAM shows a median closest to this optimal value and a smaller interquartile range compared to the others. Similarly, for the SCARED dataset (Figure 7), BodySLAM outperforms EndoDepth and EndoSfmLearner across all metrics. As with the Hamlyn dataset, BodySLAM demonstrates lower median values and smaller interquartile ranges for Absolute relative difference, RMSE, RMSE log and Squared Relative error (first row, Figure 7). For the threshold accuracy (second row, Figure 7), BodySLAM median is closest to the optimal value of 1, with a smaller interquartile range compared to the other methods. To deepen the results we also performed a statistical analysis presented in Table 6 for the Hamlyn dataset and in Table 7 for the SCARED Dataset.
The statistical analysis presented on the Hamlyn dataset shows that BodySLAM outperforms the other depth estimation models, except for the threshold accuracy $(1.25)^{2}$ and $(1.25)^{3}$, where BodySLAM does not significantly outperform EndoSfmLearner. For the SCARED dataset, BodySLAM outperforms the other depth estimation models in most metrics. However, there are a few exceptions: in RMSE, BodySLAM does not statistically outperform EndoSfmLearner; in Squared Relative Error, BodySLAM does not statistically outperform both EndoDepth and EndoSfmLearner; and in Accuracy $(1.25)^{3}$, BodySLAM does not outperform EndoSfmLearner statistically.
@@ 7 Conclusions
This study presented a robust deep learning-based SLAM approach that can effectively operate across various endoscopic surgical settings, including laparoscopy, gastroscopy, and colonoscopy. We addressed the challenges posed by hardware limitations and environmental variations by integrating deep learning models with strong generalization capabilities into the framework. Our results demonstrate that Zoe’s exceptional generalization abilities, previously showcased in various imaging contexts, can be effectively adapted to the endoscopic field. When compared to current state-of-the-art depth estimation algorithms in endoscopy, Zoe exhibited superior performance, confirming its potential for SLAM applications in surgical environments. These findings are consistent with the results reported by [43]. Another key contribution of our work is the development of a novel unsupervised pose estimation method, CyclePose, based on the CycleGAN architecture. Our comprehensive evaluations revealed CyclePose’s robust generalization capabilities and competitive performance compared to existing state-of-the-art pose estimation techniques, while maintaining the lowest inference time. This makes CyclePose a promising solution for real-time surgical applications. Furthermore, an ablation study conducted on the pose estimation module highlighted the effectiveness of new terms introduced for cycle loss, with the model incorporating the custom term consistently outperforming its counterpart without the term. These considerations are also confirmed by the statistical analysis revealing that Zoe significantly outperformed other depth estimation models in most metrics, with only a few exceptions where the differences were not statistically significant. It also showed that the CyclePose approach performed comparably to other methods while having the lowest inference time, making it a promising solution for real-time surgical applications.
@@ Acknowledgements
This work was partially founded by: i) Piano Nazionale Ripresa e Resilienza (PNRR) - HEAL ITALIA Extended Partnership - SPOKE 2 Cascade Call - "Intelligent Health" with the project BISTOURY - 3D-guided roBotIc Surgery based on advanced navigaTiOn systems and aUgmented viRtual realitY (CUP: J33C22002920006); ii) UCBM University Strategic Projects 2023 with the Proof of Concept (PoC) project BONE - Cooperative Robotic System for spinal surgery; iii) PNRR MUR project PE0000013-FAIR.

### v2

@@ 6.2 Monocular Depth Estimation Module (MDEM)
Figure 6, 7 present the performances of the monocular depth estimation module, BodySLAM, compared to EndoSfmLearner, EndoDepth and AFSfMLearner respectively on the Hamlyn and SCARED datasets.
@@
Figure 6: Comparison of depth estimation performance for EndoDepth, EndoSfmLearner, AFSfMlearner and BodySLAM across various metrics on Hamlyn Dataset. Box plots show that BodySLAM achieves the best performance with the lowest median values for absolute relative difference, RMSE, RMSE log, and squared relative difference, and the highest median values for accuracy thresholds $\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$. EndoSfmLearner demonstrates the second best performance, while EndoDepth shows the poorest results among the three methods compared.
@@
Figure 7: Comparison of depth estimation performance for EndoDepth, EndoSfmLearner, AFSfMlearner and BodySLAM across various metrics on Hamlyn Dataset. Box plots show that BodySLAM achieves the best performance with the lowest median values for absolute relative difference, RMSE log, and squared relative difference, and the highest median values for accuracy thresholds $\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$. EndoSfmLearner demonstrates the second best performance, while EndoDepth shows the poorest results among the three methods compared. Outliers beyond 1.5 times the interquartile range from the quartiles have been excluded to enhance clarity and focus on the central trends of the data.
The figures compare the three methods across multiple evaluation metrics for each dataset. The x-axis represents the different methods (EndoSfmLearner, EndoDepth, AFSfMLearner and BodySLAM), while the y-axis shows the values for the evaluation metrics.
It’s worth noting that OneSLAM is not included in this comparison. This is because OneSLAM produces sparse depth maps, while the other methods generate dense depth maps. The evaluation metrics used in this study are designed for dense depth maps, making a direct comparison with OneSLAM’s sparse output impossible. When interpreting these results, it’s important to consider the training backgrounds of the methods. EndoDepth was trained on the Hamlyn dataset, while AFSfMLearner was trained on the SCARED dataset. This prior exposure to the test data or related datasets may introduce a potential bias in their performance, which should be taken into account when interpreting their results. In contrast, BodySLAM has not been exposed to either dataset during its training, which highlights its generalization capabilities.
@@
In the case of the Hamlyn dataset (Figure 6), BodySLAM outperforms all competitors across all metrics. For Absolute relative difference, RMSE, RMSE log and Squared Relative error (first row, Figure 6), BodySLAM exhibits lower median values and smaller interquartile ranges compared to the other methods. Regarding the threshold accuracy (second row, Figure 6), where values close to or equal to 1 are considered optimal, BodySLAM shows a median closest to this optimal value and a smaller interquartile range compared to the others. Similarly, for the SCARED dataset (Figure 7), BodySLAM outperforms all competitors across all metrics. As with the Hamlyn dataset, BodySLAM demonstrates lower median values and smaller interquartile ranges for Absolute relative difference, RMSE, RMSE log and Squared Relative error (first row, Figure 7). For the threshold accuracy (second row, Figure 7), BodySLAM median is closest to the optimal value of 1, with a smaller interquartile range compared to the other methods. To deepen the results we also performed a statistical analysis presented in Table 6 for the Hamlyn dataset and in Table 7 for the SCARED Dataset. The statistical analysis presented on the Hamlyn dataset shows that BodySLAM outperforms the other depth estimation models, except for the threshold accuracy $(1.25)^{2}$ and $(1.25)^{3}$, where BodySLAM does not significantly outperform EndoSfmLearner. For the SCARED dataset, BodySLAM outperforms the other depth estimation models in most metrics. However, there are a few exceptions: in RMSE, BodySLAM does not statistically outperform EndoSfmLearner; in Squared Relative Error, BodySLAM does not statistically outperform both EndoDepth and EndoSfmLearner; and in Accuracy $(1.25)^{3}$, BodySLAM does not statistically outperform EndoSfmLearner.
@@ 7 Conclusions
This study presented a robust deep learning-based SLAM approach that can effectively operate across various endoscopic surgical settings, including laparoscopy, gastroscopy, and colonoscopy. We addressed the challenges posed by hardware limitations and environmental variations by integrating deep learning models with strong generalization capabilities into the framework. Our results demonstrate that our novel unsupervised pose estimation method, CycleVO, exhibits robust generalization capabilities and competitive performance compared to existing state-of-the-art pose estimation techniques, while maintaining the lowest inference time. This makes CycleVO a promising solution for real-time surgical applications. An ablation study conducted on the pose estimation module highlighted the effectiveness of new terms introduced for cycle loss, with the model incorporating the custom term consistently outperforming its counterpart without the term. These findings were confirmed by statistical analysis revealing that the CycleVO approach performed comparably to other methods while having the lowest inference time. Furthermore, we demonstrated that Zoe’s exceptional generalization abilities, previously showcased in various imaging contexts, can be effectively adapted to the endoscopic field. When compared to current state-of-the-art depth estimation algorithms in endoscopy, Zoe exhibited superior performance, confirming its potential for SLAM applications in surgical environments. These findings are consistent with the results reported by [44]. Statistical analysis revealed that Zoe significantly outperformed other depth estimation models in most metrics, with only a few exceptions where the differences were not statistically significant.
@@ Acknowledgements
This work was partially founded by: i) Piano Nazionale Ripresa e Resilienza (PNRR) - HEAL ITALIA Extended Partnership - SPOKE 2 Cascade Call - "Intelligent Health" with the project BISTOURY - 3D-guided roBotIc Surgery based on advanced navigaTiOn systems and aUgmented viRtual realitY (CUP: J33C22002920006); ii) UCBM University Strategic Projects 2023 with the Proof of Concept (PoC) project BONE - Cooperative Robotic System for spinal surgery; iii) PNRR MUR project PE0000013-FAIR. Resources are provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at Alvis @ C3S.